OpenStudy (anonymous):

Consider the triangle shown below find the value of sin cos tan sec csc & cot

OpenStudy (alfie):

I cannot see the picture, attach it! :)

OpenStudy (alfie):

tan = y/x = 5/12. So 0(angle) = arctan(5/12) => 22,619 You might wanna round it to 23° So you have the angle now, you can find everything easily with your calc :)

OpenStudy (anonymous):

so i'd put in sin(23)

OpenStudy (anonymous):

cause im putting that and im not getting any of the choices

OpenStudy (alfie):

Just open up windows calc, type in 23 and click on "sin" You'll get around 0,40, which is pretty reasonable.